# Recent Advances in Discrete Speech Tokens: A Review

### v2

@@ Abstract
Index Terms:
Discrete speech tokens, neural audio codec, speech tokenizer, speech LLMs, spoken language modeling, speech generation, acoustic tokens, semantic tokens
@@ I Introduction
The rapid advancement of large language models (LLMs) in natural language processing has revolutionized speech generation tasks [1, 2], with speech being tokenized and modeled using decoder-only Transformers [3]. Efforts starting from GSLM [4] and AudioLM [5] aim to develop text-free spoken LLMs, akin to how current LLM-powered chatbots enable text-based interactions. Other works, including VALL-E [6] and VioLA [7], extend this approach to conditional speech generation tasks, such as zero-shot text-to-speech and speech translation. However, this paradigm requires data to be tokenized, as LLMs typically process discrete data only. Textual tokens naturally meet this requirement because they are designed as discrete units separated by clear boundaries, whereas raw speech signals are continuous and boundary-less. Therefore, a necessary step before applying speech data to LLM is the tokenization of speech, whose goal is:
To convert long speech waveforms into short discrete tokens for downstream tasks. These tokens should be compatible with the underlying textual representations, especially for language modeling approaches targeted at speech.
@@
As a result, significant efforts have been directed towards developing efficient and powerful speech tokenization methods. Generally, these methods are based on two distinct principles, giving rise to two types of speech tokens: acoustic tokens and semantic tokens. Acoustic tokens are derived from neural codecs designed to encode speech at a low bitrate while preserving as much information as possible. In contrast, semantic tokens originate from speech self-supervised learning (SSL) [8], which aims to learn a more phonetic or semantic representation space, making it easier for speech recognition. These two nearly independent lines of research magically intersect in the context of language modeling for speech. Now, there are also efforts that try to design a speech tokenizer that accomplishes the two objectives simultaneously [9, 10]. Consequently, speech tokenization has become a core problem of speech processing under the new paradigm, with versatile downstream applications, as shown in Fig.1.
Despite the rapid development and numerous recent works, a comprehensive taxonomy of methodologies in discrete speech tokens has not been clearly constructed. Existing reviews [8, 11, 12, 1, 13, 2] in this field overlook the diverse categories and methodologies in both acoustic and semantic tokens. For example, [1, 2] focus primarily on methods in spoken language modeling, providing only brief descriptions of some speech tokens used in existing models. The taxonomy of neural audio codecs has been summarized in [12, 14], but the realm of semantic tokens is still overlooked. In this review, we provide a comprehensive overview of the concepts, methods, and characteristics of various types of discrete speech tokens, with their applications in spoken language understanding, speech generation, and spoken dialogue models. We hope that through this review, the community can have a clear understanding of the current development and key technologies of discrete speech tokens, so as to promote further research in the future.
@@
The structure of this review is shown in Fig.2. Following [5, 15, 16], we classify discrete speech tokens into acoustic and semantic tokens based on their principles. We will characterize the two types of tokens both by conceptual descriptions and unified experimental comparisons.
@@
(Section II), color=turquoise, fill=turquoise!15, text width=6cm, text=black [Offline Clustering: k-means clustering, agglomerative clustering, etc., color=turquoise, fill=turquoise!15, text width=8cm, text=black] [Vector Quantization: k-means VQ, Gumbel VQ, RVQ, GVQ, FSQ, etc., color=turquoise, fill=turquoise!15, text width=8cm, text=black ] ] [Acoustic Tokens
(Section III, Table I), color=limegreen, fill=limegreen!10, text width=2.5cm, text=black [Model architectures
(Section III-A), color=limegreen, fill=limegreen!10, text width=2.5cm, text=black [VQ-GAN: CNN-based, Transformer-based, U-Net-based, etc. , color=limegreen, fill=limegreen!10, text width=8.5cm, text=black], [Latent diffusion: LaDiffCodec [17], SemantiCodec [18], etc. , color=limegreen, fill=limegreen!10, text width=8.5cm, text=black] ], [General-Purpose
(Section III-B), color=limegreen, fill=limegreen!10, text width=2.5cm, text=black [ Advanced VQ methods and model architectures: DAC [19], TS3-Codec [20], etc. , color=limegreen, fill=limegreen!10, text width=8.5cm, text=black], [ Temporal redundancy reduction: Disen-TF-Codec [21], TiCodec [22], etc. , color=limegreen, fill=limegreen!10, text width=8.5cm, text=black], [ Multi-resolution or variable-bitrate: SNAC [23], VRVQ [24], etc. , color=limegreen, fill=limegreen!10, text width=8.5cm, text=black] ], [Semantic Distillation
(Section III-C), color=limegreen, fill=limegreen!10, text width=2.5cm, text=black [ SSL feature guidance: SpeechTokenizer [9], Mimi [10], etc. , color=limegreen, fill=limegreen!10, text width=8.5cm, text=black], [ Fixed semantic codebook: LLM-Codec [25], etc. , color=limegreen, fill=limegreen!10, text width=8.5cm, text=black], [ SSL as inputs or outputs: X-Codec [26], SemantiCodec [18], etc. , color=limegreen, fill=limegreen!10, text width=8.5cm, text=black], ], [Disentanglement
(Section III-D), color=limegreen, fill=limegreen!10, text width=2.5cm, text=black [ Gradient reversal layer: SSVC [27], FACodec [28], etc. , color=limegreen, fill=limegreen!10, text width=8.5cm, text=black], [ Perturbation: LSCodec [29], etc. , color=limegreen, fill=limegreen!10, text width=8.5cm, text=black], [ Source separation: SD-Codec [30], etc. , color=limegreen, fill=limegreen!10, text width=8.5cm, text=black], ], ] [Semantic Tokens
(Section IV, Table II), color=plum, fill=plum!15, text width=2.5cm, text=black [General-Purpose SSL
(Section IV-A), color=plum, fill=plum!15, text width=3cm, text=black [Contrastive: vq-wav2vec [31], wav2vec 2.0 [32], etc., color=plum, fill=plum!15, text width=8cm, text=black], [Predictive: HuBERT [33], WavLM [34], etc., color=plum, fill=plum!15, text width=8cm, text=black], ], [Perturbation-Invariant SSL
(Section IV-B), color=plum, fill=plum!15, text width=3.5cm, text=black [Contrastive-based: ContentVec [35], SPIRAL [36], etc., color=plum, fill=plum!15, text width=7.5cm, text=black], [Distribution-based: Spin [37], NAST [38], etc., color=plum, fill=plum!15, text width=7.5cm, text=black], [CTC-based: Gat et al. [39], etc., color=plum, fill=plum!15, text width=7.5cm, text=black], ], [Supervised Tokens
(Section IV-C), color=plum, fill=plum!15, text width=3cm, text=black [ ASR-based: Whisper [40], $\mathcal{S}^{3}$ Tokenizer [41], etc. , color=plum, fill=plum!15, text width=5.5cm, text=black], ], [    Length Reduction by Deduplication and Acoustic BPE (Section IV-D), color=plum, fill=plum!15, text=black, align=center, text width=8.5cm ], [       Variable Frame Rate Tokens and Unit Discovery (Section IV-E), color=plum, fill=plum!15, text width=8.5cm, text=black, align=center ], [                      Speech Token Vocoders (Section IV-F), color=plum, fill=plum!15, text width=8.5cm, text=black, align=center ], ], [Analysis (Section V): Metrics, benchmarks, reconstruction and voice conversion comparisons, color=lightyellow, fill=lightyellow!15, text width=11cm, text=black, align=center ], [Applications (Section VI), Challenges and Future Directions (Section VII), color=lightyellow, fill=lightyellow!15, text width=11cm, text=black, align=center ] ]
Figure 2: Structure of this review
@@ II Pre-requisites: Discrete Representation Learning
Denote $\bm{x}\in\mathbb{R}^{d}$ as a vector in the $d$-dimensional continuous space. A quantization process $q$ transforms $\bm{x}$ into a discrete token in a finite set, i.e. $q(\bm{x}):\mathbb{R}^{d}\to\{1,2,...,V\}$ where $V$ is the vocabulary size. The output tokens are sometimes referred to as indexes in the finite $V$-cardinal set. The function $q$ is usually associated with a codebook $\mathcal{C}=\{\bm{c}_{1},\bm{c}_{2},...,\bm{c}_{V}\}$ where every code-vector $\bm{c}_{i}\in\mathbb{R}^{d}$ corresponds to the $i$-th token. The code-vectors are representations of tokens in the original $d$-dimensional space. As $V$ elements can be encoded using $\lceil\log_{2}V\rceil$ raw bits [^1] We use $\lceil z\rceil$ to denote the ceiling of a scalar $z$, i.e., the smallest integer greater than or equal to $z$. Similarly, $\lfloor z\rfloor$ denotes the floor of $z$, i.e., the largest integer less than or equal to $z$., quantization often compresses the cost for data storage and transmission to a great extent.
@@ II-A Offline Clustering
Clustering is a simple approach for quantization. Given a dataset $X=\{\bm{x}_{1},\bm{x}_{2},...\bm{x}_{N}\}$, a clustering algorithm aims to assign each sample $\bm{x}_{i}$ to a group such that some cost is minimized. The most frequently used clustering method for discrete speech tokens is k-means clustering [42], e.g. in GSLM [4]. K-means is a clustering algorithm based on Euclidean distances. Its training process iteratively assigns each data sample to the nearest centroid, and moves cluster centroids till convergence, with a pre-defined number of clusters. After training, the centroids form the codebook, and new data can be quantized to the index of the nearest centroid in this Voronoi partition. In practice, centroids are usually initialized with the k-means++ algorithm [43] for better convergence.
Hierarchical agglomerative clustering has also been used in discrete speech tokens, which iteratively merges the closest clusters. It is usually applied after k-means to reduce the number of clusters [44, 45]. Other clustering algorithms are less explored in the context of discrete speech tokens.
@@ II-B Vector Quantization
Clustering is often an isolate process, thus cannot be optimized together with other neural network modules. Instead, vector quantization (VQ) [46] enables a learnable network module that allows gradients to pass through when producing discrete representations. Autoencoders with a VQ module is termed VQ-VAE [47]. There are multiple VQ methods:
@@ II-B1 K-means VQ
Then, code-vector $\bm{c}_{k}\triangleq\bm{c}_{q(\mathbf{x})}$ is fed to subsequent networks. As the $\min$ operation is not differentiable, straight-through estimators (STEs) [48] are usually applied to graft gradients, i.e. $\operatorname{STE}(\bm{c}_{k},\bm{x})=\bm{x}+\operatorname{sg}(\bm{c}_{k}-\bm{x})$ where $\operatorname{sg(\cdot)}$ stops tracking gradients. In this way, the input value to subsequent networks is still $\bm{c}_{k}$, but gradients are grafted to $\bm{x}$ in back propagation.
Auxiliary loss functions are often used together with k-means VQ [47]: commitment loss $\mathcal{L}_{\text{cmt}}=\|\operatorname{sg}(\bm{c}_{k})-\bm{x}\|^{2}$ and codebook loss $\mathcal{L}_{\text{code}}=\|\operatorname{sg}(\bm{x})-\bm{c}_{k}\|^{2}$. The commitment loss pushes the continuous input $\bm{x}$ towards the closest codebook entry, while the codebook loss does the opposite and updates the code-vector $\bm{c}_{k}$. The two loss terms are weighted by different factors to put different optimization strengths on $\bm{x}$ and $\bm{c}_{k}$, as pushing $\bm{c}_{k}$ towards $\bm{x}$ is an easier task. It is also common to replace $\mathcal{L}_{\text{code}}$ with exponential moving average (EMA) to update the codebook instead [49], which does not rely on explicit loss functions.
VQ in high-dimensional spaces is known to suffer from codebook collapse, where the codebook usage is highly imbalanced [50, 51]. To improve the utilization of codebook, random replacement (as known as codebook expiration) can be applied [51] on code-vectors that have remained inactive for a long time. Other solutions include additional auxiliary constraints such as entropy penalty [52, 53], factorized codebook lookup in low-dimensional space [54], and adding a linear projection to update all code-vectors together [55].
@@ II-B2 Gumbel VQ
Instead of quantizing by Euclidean distance, another choice is by probability. Gumbel VQ [56] uses Gumbel-Softmax as a proxy distribution for traditional Softmax to allow differentiable sampling. Given input $\bm{x}$ and a codebook of size $V$, a transform $h(\cdot)$ is applied on $\bm{x}$ into $V$ logits: $\bm{l}=h(\bm{x})\in\mathbb{R}^{V}$. In inference, quantization is performed by choosing the index with the largest logit, i.e. $q(\bm{x})=\arg\max_{i}\left\{\bm{l}^{(i)}\right\}$. In training, samples are drawn from the categorical distribution implied by $\bm{l}$ for the subsequent neural networks. To achieve efficient sampling and let gradients pass through, Gumbel trick is used:
@@
where Eq.(2) samples Gumbel noise $\bm{v}$ element-wise, and Eq.(3) calculates Gumbel-Softmax distribution $\bm{s}$ with a temperature $\tau$. The forward pass simply use $j=\arg\max_{i}\{\bm{s}^{(i)}\}$ as the sampled index, but the true gradient of Gumbel-Softmax is used in backward pass. In other words, the gradient on the one-hot distribution corresponding to $j$ is grafted to $\bm{s}$ as an approximate. The temperature $\tau$ balances the approximation accuracy and gradient variances. The transform $h(\cdot)$ is usually parameterized as neural networks, or negatively proportional to Euclidean distances [57].
@@
II-B3 Finite Scalar Quantization
As mentioned before, VQ methods based on code-vector assignment usually suffer from codebook collapse. Despite many efforts, this remains a crucial challenge. Finite scalar quantization (FSQ) [58] is an alternative to perform quantization in scalar domain. FSQ quantizes each dimension of a vector $\bm{x}$ into $L$ levels. For the $i$-th dimension $\bm{x}^{(i)}$, FSQ transforms the values to into limited range and then rounds to integers, i.e.

### v3

@@ I Introduction
The rapid advancement of large language models (LLMs) in natural language processing has revolutionized speech generation tasks [cui2024recent, ji2024wavchat], with speech being tokenized and modeled using decoder-only Transformers [transformer]. Efforts starting from GSLM [lakhotia2021generative] and AudioLM [borsos2023audiolm] aim to develop text-free spoken LLMs, akin to how current LLM-powered chatbots enable text-based interactions. Other works, including VALL-E [valle] and VioLA [wang2024viola], extend this approach to conditional speech generation tasks, such as zero-shot text-to-speech and speech translation. However, this paradigm requires data to be tokenized, as LLMs typically process discrete data only. Textual tokens naturally meet this requirement because they are designed as discrete units separated by clear boundaries, whereas raw speech signals are continuous and boundary-less. Therefore, a necessary step before applying speech data to LLM is the tokenization of speech, whose goal is:
To transform long speech waveforms into compact sequences of discrete tokens compatible with textual representations, particularly for language modeling tasks involving speech.
@@
As a result, significant efforts have been directed towards developing efficient and powerful speech tokenization methods. Generally, these methods are based on two distinct principles, giving rise to two types of speech tokens: acoustic tokens and semantic tokens. Acoustic tokens are derived from neural codecs designed to encode speech at a low bitrate while preserving as much information as possible [wu2024towards, kim2024neural, anees2024speech, du2025codecfake]. In contrast, semantic tokens originate from speech self-supervised learning (SSL) [mohamed2022self], which aims to learn a more phonetic or semantic representation space, making it easier for speech recognition. These two nearly independent lines of research magically intersect in the context of language modeling for speech. Now, there are also efforts that try to design a speech tokenizer that accomplishes the two objectives simultaneously [zhang2024speechtokenizer, kyutai2024moshi]. Consequently, speech tokenization has become a core problem of speech processing under the new paradigm, with versatile downstream applications, as shown in Fig.1.
Prior to discrete tokens, continuous speech representations from autoencoders and self-supervised learning have been extensively explored [mohamed2022self]. Comparatively, continuous representations generally provide higher fidelity in reconstruction or stronger performance on understanding tasks, but lack the compactness, symbolic abstraction, and compatibility with language model-based generation that discrete tokens easily afford. In other words, they highlight different tradeoffs in representation learning.
Despite the rapid development and numerous recent works, a comprehensive taxonomy of methodologies in discrete speech tokens has not been clearly constructed. Existing reviews [mohamed2022self, anees2024speech, wu2024towards, cui2024recent, kim2024neural, ji2024wavchat] in this field overlook the diverse categories and methodologies in both acoustic and semantic tokens. For example, [cui2024recent, ji2024wavchat] focus primarily on methods in spoken language modeling, providing only brief descriptions of some speech tokens used in existing models. The taxonomy of neural audio codecs has been summarized in [wu2024towards, du2025codecfake], but the realm of semantic tokens is still overlooked. In this review, we provide a comprehensive overview of the concepts, methods, and characteristics of various types of discrete speech tokens, with their applications in spoken language understanding, speech generation, and spoken dialogue models. We hope that through this review, the community can have a clear understanding of the current development and key technologies of discrete speech tokens, so as to promote further research in the future.
@@
The structure of this review is shown in Fig.2. Following [borsos2023audiolm, kharitonov2023speak, yang2024towards], we classify discrete speech tokens into acoustic and semantic tokens based on their principles. We will characterize the two types of tokens both by conceptual descriptions and unified experimental comparisons.
@@
[…]
Discrete Speech Tokens Pre-requisites: Discrete Representation Learning (Section II) Offline Clustering: k-means clustering, agglomerative clustering, etc. Vector Quantization: k-means VQ, Gumbel VQ, FSQ, GVQ, RVQ, etc. Acoustic Tokens (Section III, Table II) Model architectures (Section III-A) VQ-GAN: CNN-based, Transformer-based, U-Net-based, etc. Diffusion: LaDiffCodec [yang2024generative], SemantiCodec [liu2024semanticodec], etc. General-Purpose (Section III-B) Advanced VQ methods and model architectures: DAC [kumar2024high], TS3-Codec [wu2024ts3codectransformerbasedsimplestreaming], etc. Temporal redundancy reduction: Disen-TF-Codec [jiang2023disentangled], TiCodec [ticodec], etc. Multi-resolution or variable-bitrate: SNAC [Siuzdak_SNAC_Multi-Scale_Neural_2024], VRVQ [chae2024variable], etc. Semantic Distillation (Section III-C) Semantic feature guidance: SpeechTokenizer [zhang2024speechtokenizer], Mimi [kyutai2024moshi], etc. Fixed semantic codebook: LLM-Codec [yang2024uniaudio15], etc. Semantic features as inputs or outputs: X-Codec [ye2024codec], SemantiCodec [liu2024semanticodec], etc. Disentanglement (Section III-D) Gradient reversal layer: SSVC [SSVC], FACodec [facodec], etc. Perturbation: LSCodec [guo2024lscodec], etc. Source separation: SD-Codec [bie2024learning], DeCodec [luo2025decodec], etc. Semantic Tokens (Section IV, Table III) SSL Semantic Tokens (Section IV-A) Contrastive models: vq-wav2vec [vq-wav2vec], wav2vec 2.0 [baevski2020wav2vec], etc. Predictive models: HuBERT [hsu2021hubert], WavLM [chen2022wavlm], etc. Perturbation-invariant SSL models: ContentVec [qian2022contentvec], NAST [messica2024nast], etc. Supervised Semantic Tokens (Section IV-B): Whisper [whisper], $\mathcal{S}^{3}$ Tokenizer [du2024cosyvoice], etc. Speech Token Vocoders (Section IV-C) Length Reduction and Variable-Rate Tokenization (Section V)     Length Reduction by Deduplication and Acoustic BPE (Section V-A)        Variable Frame Rate Tokens and Unit Discovery (Section V-B) Analysis (Section VI): Metrics, benchmarks, experimental comparisons (reconstruction, voice conversion, downstream semantic modeling) Applications (Section VII), Challenges and Future Directions (Section VIII)
Figure 2: Structure of this review. After a brief introduction to the preliminary knowledge, we will taxonomize acoustic tokens and semantic tokens, followed by cross-cutting methods such as length reduction and variable-rate tokenization. Later sections cover experimental analysis, applications, and future directions. Each branch corresponds to a subsection in the paper.
@@ II Pre-requisites: Discrete Representation Learning
Denote $\bm{x}\in\mathbb{R}^{d}$ as a vector in the $d$-dimensional continuous space. A quantization process $q$ transforms $\bm{x}$ into a discrete token in a finite set, i.e. $q(\bm{x}):\mathbb{R}^{d}\to\{1,2,...,V\}$ where $V$ is the vocabulary size. The output tokens are sometimes referred to as indexes in the finite $V$-cardinal set. The function $q$ is usually associated with a codebook $\mathcal{C}=\{\bm{c}_{1},\bm{c}_{2},...,\bm{c}_{V}\}$ where every code-vector $\bm{c}_{i}\in\mathbb{R}^{d}$ corresponds to the $i$-th token. The code-vectors are representations of tokens in the original $d$-dimensional space. As $V$ elements can be encoded using $\lceil\log_{2}V\rceil$ raw bits [^1] We denote by $\lceil z\rceil$ and $\lfloor z\rfloor$ the ceiling and floor of scalar $z$, i.e., $\lceil z\rceil=\min\{n\in\mathbb{Z}:n\geq z\}$ and $\lfloor z\rfloor=\max\{n\in\mathbb{Z}:n\leq z\}$. , quantization often compresses the cost for data storage and transmission to a great extent.
@@ II-A Offline Clustering
Clustering is a simple approach for quantization. Given a dataset $X=\{\bm{x}_{1},\bm{x}_{2},...\bm{x}_{N}\}$, a clustering algorithm aims to assign each sample $\bm{x}_{i}$ to a group such that some cost is minimized. The most frequently used clustering method for discrete speech tokens is k-means clustering [IKOTUN2023178]. For example, k-means is applied on HuBERT [hsu2021hubert] features in GSLM [lakhotia2021generative]. K-means is a clustering algorithm based on Euclidean distances. Its training process iteratively assigns each data sample to the nearest centroid, and moves cluster centroids till convergence, with a pre-defined number of clusters. After training, the centroids form the codebook, and new data can be quantized to the index of the nearest centroid in this Voronoi partition. In practice, centroids are usually initialized with the k-means++ algorithm [kmeans++] for better convergence.
Hierarchical agglomerative clustering has also been used in discrete speech tokens, which iteratively merges the closest clusters. It is usually applied after k-means to reduce the number of clusters [cho2024sd, baade2024syllablelm]. Other clustering algorithms are less explored in the context of discrete speech tokens.
@@ II-B Vector Quantization
Clustering is often an isolate process, thus cannot be optimized together with other neural network modules. Instead, vector quantization (VQ) [gray1984vector] enables a learnable network module that allows gradients to pass through when producing discrete representations. Autoencoders with a VQ module is termed VQ-VAE [VQVAE]. There are multiple VQ methods:
@@ II-B1 K-means VQ
Then, code-vector $\bm{c}_{k}\triangleq\bm{c}_{q(\mathbf{x})}$ is fed to subsequent networks. As the $\min$ operation is not differentiable, straight-through estimators (STEs) [bengio2013estimating] are usually applied to graft gradients, i.e. $\operatorname{STE}(\bm{c}_{k},\bm{x})=\bm{x}+\operatorname{sg}(\bm{c}_{k}-\bm{x})$ where $\operatorname{sg(\cdot)}$ stops tracking gradients. In this way, the input value to subsequent networks is still $\bm{c}_{k}$, but gradients are grafted to $\bm{x}$ in back propagation.
Auxiliary loss functions are often used together with k-means VQ [VQVAE]: commitment loss $\mathcal{L}_{\text{cmt}}=\|\operatorname{sg}(\bm{c}_{k})-\bm{x}\|^{2}$ and codebook loss $\mathcal{L}_{\text{code}}=\|\operatorname{sg}(\bm{x})-\bm{c}_{k}\|^{2}$. The commitment loss pushes the continuous input $\bm{x}$ towards the closest codebook entry, while the codebook loss does the opposite and updates the code-vector $\bm{c}_{k}$. The two loss terms are weighted by different factors to put different optimization strengths on $\bm{x}$ and $\bm{c}_{k}$, as pushing $\bm{c}_{k}$ towards $\bm{x}$ is an easier task. It is also common to replace $\mathcal{L}_{\text{code}}$ with exponential moving average (EMA) to update the codebook instead [razavi2019generating], which does not rely on explicit loss functions.
VQ in high-dimensional spaces is known to suffer from codebook collapse, where the codebook usage is highly imbalanced [lancucki2020robust, dhariwal2020jukebox]. To improve codebook utilization, random replacement (as known as codebook expiration) can be applied [dhariwal2020jukebox] on code-vectors that have remained inactive for a long time. Other popular solutions include additional auxiliary constraints such as entropy penalty [chang2022maskgit, yu2024language], factorized codebook lookup in low-dimensional space [yu2022vectorquantized], and reparameterizing code-vectors through a learnable linear projection [zhu2024addressing].
@@ II-B2 Gumbel VQ
Instead of quantizing by Euclidean distance, another choice is by probability. Gumbel VQ [jang2017categorical] uses Gumbel-Softmax as a proxy distribution for traditional Softmax to allow differentiable sampling. Given input $\bm{x}$ and a codebook of size $V$, a transform $h(\cdot)$ is applied on $\bm{x}$ into $V$ logits: $\bm{l}=h(\bm{x})\in\mathbb{R}^{V}$. In inference, quantization is performed by choosing the index with the largest logit, i.e. $q(\bm{x})=\arg\max_{i}\left\{\bm{l}^{(i)}\right\}$. In training, samples are drawn from the categorical distribution implied by $\bm{l}$ for the subsequent neural networks. To achieve efficient sampling and let gradients pass through, Gumbel trick is used:
@@
where Eq.(2) samples Gumbel noise $\bm{v}$ element-wise, and Eq.(3) calculates Gumbel-Softmax distribution $\bm{s}$ with a temperature $\tau$. The forward pass simply uses $j=\arg\max_{i}\{\bm{s}^{(i)}\}$ as the sampled index, but the true gradient of Gumbel-Softmax is used in backward pass. In other words, the gradient on the one-hot distribution corresponding to $j$ is grafted to $\bm{s}$ as an approximate. The temperature $\tau$ balances the approximation accuracy and gradient variances. The transform $h(\cdot)$ is usually parameterized as neural networks, or negatively proportional to Euclidean distances [jiang2023latent].
@@
II-B3 Finite Scalar Quantization (FSQ)
As mentioned before, VQ methods based on code-vector assignment usually suffer from codebook collapse. Despite many efforts, this remains a crucial challenge. FSQ [mentzer2024finite] is an alternative that performs quantization in scalar domain. FSQ quantizes each dimension of a vector $\bm{x}$ into $L$ levels. For the $i$-th dimension $\bm{x}^{(i)}$, FSQ transforms the values into a limited range and then rounds to integers, i.e.